OpenStudy (anonymous):

Find the area of a triangle whose base is 5cm and whose height is 12cm.

OpenStudy (sleepyhead314):

Hello and Welcome to OpenStudy! :) the area of a triangle = (1/2) (base) times (height) so = (1/2) times (5) times (12) = ?

OpenStudy (anonymous):

Which would be 30!